A manager is going to purchase new processing equipment and must decide on the number of spare parts to order with the new equipment. The spares cost $170 each, and any unused spares will have an expected salvage value of $40 each. The probability of usage can be described by this distribution:

Number 0 1 2 3
Probability .05 .40 .20 .35


If a part fails and a spare is not available, 2 days will be needed to obtain a replacement and install it. The cost for idle equipment is $550 per day. What quantity of spares should be ordered?

a. Use the ratio method. (Round your answer to 2 decimal places.)

Service level is  , so we need to order  Spares.
  
b. Use the tabular method.
Answer #1

Cost of Stock out Cs = 550*2 = 1100

Cost of Excess Inventory Ce= 170 - 40 = 130

Service Level = Cs/(Cs+Ce)

Service Level = 1100/(1100+130)

Service Level = 0.8943

Below is the probability table -

Number Probability Cumulative Probability
0 0.05 0.05
1 0.4 0.45
2 0.2 0.65
3 0.35 1

0.8943 lies between, 0.65 and 1, Hence

Number of Spare = 3

2)

Tabular Method -

Below is the calculation performed as per tabular method -

Demand = 0 Demand = 1 Demand = 2 Demand = 3
Stocking Level 0.05 0.4 0.2 0.35 Expected Cost
0 0 440 440 1155 2035
1 6.5 0 220 770 996.5
2 13 52 0 385 450
3 19.5 104 26 0 149.5

Cost is lowest with 3 spare parts, hence 3 spare parts should be ordered.